# RSRNav: Reasoning Spatial Relationship for Image-Goal Navigation

### v1

@@ Abstract
keywords:
Image-goal navigation, spatial relationship reasoning, visual correlation.
journal: Pattern Recognition
1 Introduction
Visual navigation [1, 2] emerges as a pivotal research area within embodied artificial intelligence, requiring an agent to navigate unfamiliar environments to reach specific goals. Although multiple information carriers can represent goals, images offer a more accurate and detailed description for specifying the goal’s location. Our research focuses on the image-goal navigation task (ImageNav), a field brimming with diverse potential applications, such as household robots, augmented reality systems, and assistance for individuals with visual impairments.
Fig 1: Illustration of human-intuitive navigation, with action decisions guided by constantly reasoning spatial relationship between the current observation $\mathbf{I}_{1},\mathbf{I}_{2},\mathbf{I}_{3}$ and the goal (perception). Goal-irrelevant $\mathbf{I}_{1}$: the current observation is irrelevant to the goal; we should either turn around or explore a new area. Goal-relevant $\mathbf{I}_{2}$: the element relevant to the goal is encountered, i.e., the bed; we stop exploring and approach the goal under the bed’s spatial guidance. Goal-proximal $\mathbf{I}_{3}$: We are close to the goal; we precisely adjust the angle based on the relative positional relationship of the goal in the current observation.
@@ 1 Introduction
More human-intuitive navigation is achieved efficiently by reasoning about the spatial relationship between the current environment and the target position. This spatial relationship can provide rich orientation information to guide navigation actions. As shown in Fig 1, the agent explores new areas when the observation is goal-irrelevant ($\mathbf{I}_{1}$); approaches the goal when the observation is goal-relevant ($\mathbf{I}_{2}$); and precisely adjusts the angle when the observation is goal-proximal ($\mathbf{I}_{3}$). Furthermore, spatial relationship between images are less affected by inconsistent viewpoints because they focus on how images relate to each other rather than on the detailed semantics information. Inspired by these, we aim to overcome the above challenges by reasoning spatial relationship and using it to guide navigation actions.
@@
The remainder of the paper is organized as follows. We briefly review the related work in Section 2. Subsequently, we present the technical details of the proposed method in Section 3. The experimental results are presented in Section 4. Finally, we conclude this paper in Section 5.
2 Related Work
2.1 Modular Methods
Modular methods for navigation break down the high-level tasks into a series of independent sub-tasks or modules responsible for handling a specific function within the navigation process, such as classic perception [10], localization, map construction (SLAM [11]), and path planning. Some of these modules require no extra training and fine-tuning [12], and some need additional data to support the training [13, 14]. As an example, [12] tackles various sub-tasks, including exploration, goal re-identification, goal localization, and local navigation. This work uses feature-matching methods to re-identify the goal in egocentric vision and then project the matched features onto a map for localization. All of these tasks are addressed using standard components, without requiring any finetuning. Some methods build semantic maps [15, 16] to help agents with better localization and planning. For instance, [15] introduces topological models of space that incorporate semantic information, enabling coarse geometric reasoning and improving resilience to actuation errors. By storing semantic features at nodes, the system can exploit statistical patterns, facilitating efficient exploration in unfamiliar environments. Some methods discard map construction and simply use modern vision tasks, such as object detection or segmentation [13, 17].
2.2 End-to-end Methods
@@
The core of end-to-end methods is to train a perception-to-action policy. CRL [20] simultaneously trains a reinforcement learning policy and a visual representation model. The policy is designed to maximize the errors in the representation model, which encourages the agent to explore its environment. As the policy generates increasingly challenging data, the representation model becomes progressively more refined and robust. ZER [5] introduces a zero-shot transfer learning framework that incorporates an innovative reward mechanism for its semantic search policy. This approach allows the model to effectively transfer knowledge gained from a source task to multiple target tasks (such as ObjectNav, RoomNav, and ViewNav), which involve diverse goal modalities like images, sketches, audio, and labels. Additionally, this model supports zero-shot experience learning, enabling it to address target tasks without requiring any task-specific interactive training. Some other efforts and advances have focused on enhancing the agent’s perceptual capabilities, i.e., how to encode the observation’s features and the goal better. For example, ZSON [7] uses the CLIP model to improve image embeddings, which has been widely used in the multimodal field [21, 22]. OVRL [4] introduces a two-step approach: (1) offline pretraining of visual features through self-supervised learning (SSL) on a large dataset of pre-rendered indoor environment images, and (2) online fine-tuning of visuomotor representations for specific tasks, utilizing image augmentations and extended training periods. After that, OVRL-V2 [6] introduces a stronger model, pre-training vision transformers (ViT). FGPrompt [23] proposes fusion strategies to extract the features of observation based on the goal prompt. And the memory module has been introduced to address the long-horizon planning [1]. In distinction to previous works focusing on extracting better and more efficient features, we propose to model the spatial relationship between current observation and the goal and train a perception-relationship-action policy.
2.3 Visual Correlation
Identifying correlations between an image pair is critical to many computer vision tasks. For example, the image alignment task accomplishes image-level alignment ((Geometric [24] or Semantic [25, 26] alignment)) by constructing correlations between images; the object tracking task [27, 28] and multi-object tracking task [29, 30] perform instance-level association through correlation between consecutive video frames; correlation assists the optical flow task [31, 32] in obtaining pixel-level offsets. In visual navigation, [33] uses the correlation between two consecutive observations to monitor whether it is traveling on a predetermined route. We innovatively construct the visual correlation between the goal and egocentric images to model spatial relationship , guiding action prediction.
3 Method
3.1 Task Definition and Overview of RSRNav
@@
3.2 Perception
Unlike prior methods [7, 4] that rely on complex, pre-trained encoder networks, our method employs a simple ResNet-9 without any pre-training. Furthermore, we utilize a weight-sharing network for both rather than using separate networks to extract features from egocentric and goal images. At time step $t$, we employ a weight-sharing ResNet-9 to encode both the goal image $\mathbf{I}_{g}$ and the current egocentric image $\mathbf{I}_{t}$ into features (feature vectors or feature maps), as shown in Fig 2.
Fig 2: Overview of our RSRNav. “Minimalist Relationship” constructs sparse correlations to initially explore the importance of spatial relationship in ImageNav. “Dense Relationship” uses cross-correlation with finer granularity to establish richer and more detailed relationship. “Direction-aware Relationship” introduces direction information to support the agent in moving more efficiently and adjusting the angle more precisely.
3.3 Reasoning Spatial Relationship
3.3.1 Minimalist Realtionship
@@ 3.3.2 Dense Realtionship
Fig 3: Graphical representation of the correlation pyramid and center-based direction-aware searching region. The black grids represent the approximate receptive fields of the feature points of $\mathbf{F}_{t}$ and the regions of $\mathbf{F}_{g}$ at different scales on the raw image in the correlation pyramid. The light blue areas depict the searching region template for each feature point in $\mathbf{F}_{t}$, covering the surrounding area in all directions.
3.3.3 Direction-aware Realtionship
@@
where $s$ is the layer of pyramid and $\mathbf{C}_{0}$ is initialized as $\mathbf{C}$. This set of matrices provides the similarity between each feature point in $\mathbf{F}_{t}$ and every region in $\mathbf{F}_{g}$ at different scales. The black grids that we illustrate in Fig 3 represent the approximate receptive fields of the feature points in $\mathbf{F}_{t}$ and the regions in $\mathbf{F}_{g}$ at different scales on the raw image.
@@
$\mathcal{P}\left(\mathbf{x}\right)_{\emph{s}}$ defines a square region of size $3\times 3$ for each position $\mathbf{x}$ in $\mathbf{F}_{t}$. Cropping this region in the channel corresponding to $\mathbf{x}$ in the correlation matric $\mathbf{C}_{s}$, followed by reshaping, results in a $3^{2}$-dimensional correlation vector. In Fig 3, the correlation vectors for each scale are depicted within the light blue regions of the goal image. Collecting these vectors for all $\mathbf{x}$ in $\textbf{F}_{t}$ leads to $\mathbf{O}_{s}\in\mathbb{R}^{H\times W\times 3^{2}}$.
@@
3.4 Learning Navigation Policy
We train the navigation policy $\pi$ using reinforcement learning. As shown in Fig 2, the agent obtains its current correlation cue $\mathbf{C}^{\text{cue}}$ and passes it to the policy $\pi$. Following previous methods [7, 23, 5], the policy further encodes this information along with the history so far to produce a state embedding $\mathbf{s}_{t}$. An actor-critic network leverages $\mathbf{s}_{t}$ to predict state value $\mathbf{c}_{t}$ and the agent’s next action $\mathbf{a}_{t}$. The model is trained end-to-end using PPO [34] with the reward in [5].
@@ 3.4 Learning Navigation Policy
4 Experiments
4.1 Setting
Task Setup and Evaluation Metrics.    We adopt the image-goal navigation task defined in Section 3.1. Navigation is considered successful when the agent stops within 1 m of $\mathbf{p}_{g}$, within the maximum of 500 steps per episode. Following [5, 4, 23], the action space of the agent is {move forward 0.25m, turn left $30^{\circ}$, turn right $30^{\circ}$, stop}. The agent uses only an egocentric RGB camera of 128 × 128 resolution and a $90^{\circ}$ FoV sensor. To quantitatively evaluate navigation performance, we employ the success rate (SR) and Success weighted by Path Length (SPL), which takes into account navigation’s path efficiency [35].
Goal Camera Setting.    We provide evaluation results under two goal camera settings, as described in Section 3.1. Under the setting of “agent-matched goal”, the parameters of the goal camera are consistent with the agent’s egocentric camera. For the “user-matched goal” setting, the goal camera parameters are more closely aligned with the human user. We sample the height of goal camera from $\mathcal{U}(0.8m,1.5m)$, pitch delta from $\mathcal{U}(-5^{\circ},5^{\circ})$, and the HFOV from $\mathcal{U}(60^{\circ},120^{\circ})$ following [23, 36].
4.2 Comparison with State-of-the-art Methods
To enable the evaluation closer to the user experience, we add a new setting “user-matched goal” in addition to the original “agent-matched goal” setting (as described in Section 4.1). We compare the results under these two settings for each of the three datasets.
Table 1: Comparison with the state-of-the-art methods on Gibson. The two settings, “agent-matched goal” and “user-matched goal”, represent the goal camera reflecting the task executor and task publisher, respectively. The best results for each metric are highlighted in bolded. We report the ImageNav results averaged over three random seeds.
@@ 4.2 Comparison with State-of-the-art Methods
Table 2: Comparison of cross-domain generalization performance on MP3D. All methods are trained on Gibson and evaluated on MP3D. The two settings, “agent-matched goal” and “user-matched goal”, represent the goal camera reflecting the task executor and task publisher, respectively. The best results for each metric are highlighted in bolded. We report the ImageNav results averaged over three random seeds.
@@
Table 3: Comparison of cross-domain generalization performance on HM3D. All methods are trained on Gibson and evaluated on HM3D. All methods are trained on Gibson and evaluated on MP3D. The two settings, “agent-matched goal” and “user-matched goal”, represent the goal camera reflecting the task executor and task publisher, respectively. The best results for each metric are highlighted in bolded. We report the ImageNav results averaged over three random seeds.
@@
Results on Gibson.    For the “agent-matched goal” setting, as shown in Table 1, our RSRNav outperforms the state-of-the-art methods, i.e., ranks first for metrics SPL, SR. Our baseline (ZER) achieves 29.2% success, with our RSRNav outperforming it 3-fold. Meanwhile, it achieves 21.6% SPL, and our RSRNav outperforms it by 3.2x. The significant improvement demonstrates the effectiveness of our correlation-based method in ImageNav. Furthermore, our RSRNav shows considerable performance improvements compared to OVRL-V2, which uses a strong visual backbone (ViT) with pre-training. Although employing a simpler backbone, ResNet9, without pre-training, we still perform better than it in both SPL (69.5% vs. 58.7%) and SR (91.1% vs. 82.0%). This suggests that correlation information is more critical for predicting navigation actions than the rich semantic features extracted by complex encoders. Compared to Fgprompt, both using ResNet9, we surpass it by 3.0 for SPL and 0.7 for SR, the improvement in SPL demonstrates that our explicit modeling of correlation and the introduction of directional information can make navigation more efficient.
@@
Cross-Domain Generalization.    Beyond the visual domain disparity among these datasets, MP3D scenes are both more intricate and expansive compared to Gibson, while HM3D showcases a wide variety of scene types. This presents a very challenging cross-domain evaluation setting. Following the setting and testing episodes in ZER, we directly test our Gibson-trained model on these two new datasets without any tuning. As shown in Table 2, our RSRNav achieves the best results on SPL and SR under both two goal camera settings in MP3D. It is worth noting that our advantage is amplified under the “user-matched goal” setting, which again proves that our method is more generalizable and robust compared to previous methods. Similar results appear in HM3D, as shown in Table 3, although our success rate is lower than FGPrompt for setting “agent-matched goal” by 1.3, it is significantly higher in setting “user-matched goal” by 6.9.
Table 4: Comparison of our three progressive correlation-based ImageNav methods on Gibson. Baseline* represents the feature map version baseline. All ablation studies are training 100 million steps using the Gibson dataset. “MC” represents the Minimalist Correlation. “DC” represents the Dense Correlation. “Direct-C” represens the Direction-aware Correlation. All methods are trained on Gibson and evaluated on MP3D. The two settings, “agent-matched goal” and “user-matched goal”, represent the goal camera reflecting the task executor and task publisher, respectively. We report the ImageNav results averaged over three random seeds.
@@
4.3 Ablation Study
The “Navigation Via Minimalist Correlation” version is designed to explore the importance of correlation among a large pile of information, which only passes two correlation scores to the policy network. As illustrated in Table 4, it outperforms Baseline by (+2.7 SPL, +6.4 SR), indicating that despite the low information content of “Minimalist Correlation”, it is more useful and effective and that there is information redundancy in Baseline. With this exciting phenomenon, we further enhance the construction and utilization of correlation and evolve it to the “Navigation Via Dense Correlation” version. Both metrics improved by more than 3x with the richer amount of correlation information, further confirming the critical role of correlation in ImageNav. Also, with the same processing of the feature maps we bring 4× improvements in the success rate and 3× improvements in SPL comparing with Baseline*. Finally, we introduce “Direction-aware Correlation” to provide more direction information to cue ImageNav, which brings a great boost (+8.0 SPL, +6.8 SR).
Table 5: Ablation studies of “Dense Correlation” on Gibson.
@@ 4.3 Ablation Study
Table 6: Ablation studies of “Direction-aware Correlation” on Gibson.

### v2

@@ Abstract
Index Terms:
Image-goal navigation, spatial relationship reasoning, visual correlation
I Introduction
Visual navigation [1, 2, 3, 4] emerges as a pivotal research area within embodied artificial intelligence, requiring an agent to navigate unfamiliar environments to reach specific goals. Although multiple information carriers can represent goals, images offer a more accurate and detailed description for specifying the goal’s location. Our research focuses on the image-goal navigation task (ImageNav), a field brimming with diverse potential applications, such as household robots, augmented reality systems, and assistance for individuals with visual impairments.
Figure 1: Illustration of human-intuitive navigation, with action decisions guided by constantly reasoning spatial relationship between the current observation $\mathbf{I}_{1},\mathbf{I}_{2},\mathbf{I}_{3}$ and the goal (perception). Goal-irrelevant $\mathbf{I}_{1}$: the current observation is irrelevant to the goal; we should either turn around or explore a new area. Goal-relevant $\mathbf{I}_{2}$: the element relevant to the goal is encountered, i.e., the bed; we stop exploring and approach the goal under the bed’s spatial guidance. Goal-proximal $\mathbf{I}_{3}$: We are close to the goal; we precisely adjust the angle based on the relative positional relationship of the goal in the current observation.
@@ I Introduction
More human-intuitive navigation is achieved efficiently by reasoning about the spatial relationship between the current environment and the target position. This spatial relationship can provide rich orientation information to guide navigation actions. As shown in Figure 1, the agent explores new areas when the observation is goal-irrelevant ($\mathbf{I}_{1}$); approaches the goal when the observation is goal-relevant ($\mathbf{I}_{2}$); and precisely adjusts the angle when the observation is goal-proximal ($\mathbf{I}_{3}$). Furthermore, spatial relationship between images are less affected by inconsistent viewpoints because they focus on how images relate to each other rather than on the detailed semantics information. Inspired by these, we aim to overcome the above challenges by reasoning spatial relationship and using it to guide navigation actions.
@@
The remainder of the paper is organized as follows. We briefly review the related work in Section II. Subsequently, we present the technical details of the proposed method in Section III. The experimental results are presented in Section IV. Finally, we conclude this paper in Section V.
II Related Work
II-A Modular Methods
Modular methods for navigation break down the high-level tasks into a series of independent sub-tasks or modules responsible for handling a specific function within the navigation process, such as classic perception [12], localization, map construction (SLAM [13]), and path planning. Some of these modules require no extra training and fine-tuning [14], and some need additional data to support the training [15, 16]. As an example, [14] tackles various sub-tasks, including exploration, goal re-identification, goal localization, and local navigation. This work uses feature-matching methods to re-identify the goal in egocentric vision and then project the matched features onto a map for localization. All of these tasks are addressed using standard components, without requiring any finetuning. Some methods build semantic maps [17, 18] to help agents with better localization and planning. For instance, [17] introduces topological models of space that incorporate semantic information, enabling coarse geometric reasoning and improving resilience to actuation errors. By storing semantic features at nodes, the system can exploit statistical patterns, facilitating efficient exploration in unfamiliar environments. The recent work [19] uses Superglue [20] to provide directional guidance by matching the current and goal images’ keypoints. Some methods discard map construction and simply use modern vision tasks, such as object detection or segmentation [15, 21].
II-B End-to-end Methods
@@
The core of end-to-end methods is to train a perception-to-action policy. CRL [24] simultaneously trains a reinforcement learning policy and a visual representation model. The policy is designed to maximize the errors in the representation model, which encourages the agent to explore its environment. As the policy generates increasingly challenging data, the representation model becomes progressively more refined and robust. ZER [7] introduces a zero-shot transfer learning framework that incorporates an innovative reward mechanism for its semantic search policy. This approach allows the model to effectively transfer knowledge gained from a source task to multiple target tasks (such as ObjectNav, RoomNav, and ViewNav), which involve diverse goal modalities like images, sketches, audio, and labels. Additionally, this model supports zero-shot experience learning, enabling it to address target tasks without requiring any task-specific interactive training. Some other efforts and advances have focused on enhancing the agent’s perceptual capabilities, i.e., how to encode the observation’s features and the goal better. For example, ZSON [9] uses the CLIP model to improve image embeddings, which has been widely used in the multimodal field [25, 26, 27, 28]. OVRL [6] introduces a two-step approach: (1) offline pretraining of visual features through self-supervised learning (SSL) on a large dataset of pre-rendered indoor environment images, and (2) online fine-tuning of visuomotor representations for specific tasks, utilizing image augmentations and extended training periods. After that, OVRL-V2 [8] introduces a stronger model, pre-training vision transformers (ViT). FGPrompt [29] proposes fusion strategies to extract the features of observation based on the goal prompt. And the memory module has been introduced to address the long-horizon planning [1]. In distinction to previous works focusing on extracting better and more efficient features, we propose to model the spatial relationship between current observation and the goal and train a perception-relationship-action policy.
III Method
III-A Task Definition and Overview of RSRNav
@@
Figure 2: Overview of our RSRNav. “Minimalist Relationship” constructs sparse correlations to initially explore the importance of spatial relationship in ImageNav. “Dense Relationship” uses cross-correlation with finer granularity to establish richer and more detailed relationship. “Direction-aware Relationship” introduces direction information to support the agent in moving more efficiently and adjusting the angle more precisely.
@@
III-B Perception
Unlike prior methods [9, 6] that rely on complex, pre-trained encoder networks, our method employs a simple ResNet-9 without any pre-training. Furthermore, we utilize a weight-sharing network for both rather than using separate networks to extract features from egocentric and goal images. At time step $t$, we employ a weight-sharing ResNet-9 to encode both the goal image $\mathbf{I}_{g}$ and the current egocentric image $\mathbf{I}_{t}$ into features (feature vectors or feature maps), as shown in Figure 2.
III-C Reasoning Spatial Relationship
III-C1 Minimalist Realtionship
@@ III-C2 Dense Realtionship
Figure 3: Graphical representation of the correlation pyramid and center-based direction-aware searching region. The black grids represent the approximate receptive fields of the feature points of $\mathbf{F}_{t}$ and the regions of $\mathbf{F}_{g}$ at different scales on the raw image in the correlation pyramid. The light blue areas depict the searching region template for each feature point in $\mathbf{F}_{t}$, covering the surrounding area in all directions.
III-C3 Direction-aware Realtionship
@@
where $s$ is the layer of pyramid and $\mathbf{C}_{0}$ is initialized as $\mathbf{C}$. This set of matrices provides the similarity between each feature point in $\mathbf{F}_{t}$ and every region in $\mathbf{F}_{g}$ at different scales. The black grids that we illustrate in Figure 3 represent the approximate receptive fields of the feature points in $\mathbf{F}_{t}$ and the regions in $\mathbf{F}_{g}$ at different scales on the raw image.
@@
$\mathcal{P}\left(\mathbf{x}\right)_{\emph{s}}$ defines a square region of size $3\times 3$ for each position $\mathbf{x}$ in $\mathbf{F}_{t}$. Cropping this region in the channel corresponding to $\mathbf{x}$ in the correlation matric $\mathbf{C}_{s}$, followed by reshaping, results in a $3^{2}$-dimensional correlation vector. In Figure 3, the correlation vectors for each scale are depicted within the light blue regions of the goal image. Collecting these vectors for all $\mathbf{x}$ in $\textbf{F}_{t}$ leads to $\mathbf{O}_{s}\in\mathbb{R}^{H\times W\times 3^{2}}$.
@@
III-D Learning Navigation Policy
We train the navigation policy $\pi$ using reinforcement learning. As shown in Figure 2, the agent obtains its current correlation cue $\mathbf{C}^{\text{cue}}$ and passes it to the policy $\pi$. Following previous methods [9, 29, 7], the policy further encodes this information along with the history so far to produce a state embedding $\mathbf{s}_{t}$. An actor-critic network leverages $\mathbf{s}_{t}$ to predict state value $\mathbf{c}_{t}$ and the agent’s next action $\mathbf{a}_{t}$. The model is trained end-to-end using PPO [30] with the reward in [7].
@@ III-D Learning Navigation Policy
TABLE I: Comparison with the state-of-the-art methods on Gibson. The two settings, “agent-matched goal” and “user-matched goal”, represent the goal camera reflecting the task executor and task publisher, respectively. The best results for each metric are highlighted in bolded. We report the ImageNav results averaged over three random seeds.
@@
IV Experiments
IV-A Setting
Task Setup and Evaluation Metrics.   We adopt the image-goal navigation task defined in Section III-A. Navigation is considered successful when the agent stops within 1 m of $\mathbf{p}_{g}$, within the maximum of 500 steps per episode. Following [7, 6, 29], the action space of the agent is {move forward 0.25m, turn left $30^{\circ}$, turn right $30^{\circ}$, stop}. The agent uses only an egocentric RGB camera of 128 × 128 resolution and a $90^{\circ}$ FoV sensor. To quantitatively evaluate navigation performance, we employ the success rate (SR) and Success weighted by Path Length (SPL), which takes into account navigation’s path efficiency [33].
Goal Camera Setting.   We provide evaluation results under two goal camera settings, as described in Section III-A. Under the setting of “agent-matched goal”, the parameters of the goal camera are consistent with the agent’s egocentric camera. For the “user-matched goal” setting, the goal camera parameters are more closely aligned with the human user, which is inconsistent with the agent’s egocentric camera. We sample the height of goal camera from $\mathcal{U}(0.8m,1.5m)$, pitch delta from $\mathcal{U}(-5^{\circ},5^{\circ})$, and the HFOV from $\mathcal{U}(60^{\circ},120^{\circ})$ following [29, 34].
IV-B Comparison with State-of-the-art Methods
To enable the evaluation closer to the user experience, we add a new setting “user-matched goal” in addition to the original “agent-matched goal” setting (as described in Section IV-A). We compare the results under these two settings for each of the three datasets.
TABLE II: Comparison of cross-domain generalization performance on MP3D. All methods are trained on Gibson and evaluated on MP3D. The two settings, “agent-matched goal” and “user-matched goal”, represent the goal camera reflecting the task executor and task publisher, respectively.
@@ IV-B Comparison with State-of-the-art Methods
Figure 4: Comparison of convergence rates and tests results on metrics “Distance to Goal” and “Angle to Goal”. The former metric measures the distance between the agent and the goal when stopping; the latter metric is the angle between the agent pose and the goal view when stopping. We also provide the average episode test results, plotted to the right of the convergence curve.
TABLE III: Comparison of cross-domain generalization performance on HM3D. All methods are trained on Gibson and evaluated on HM3D. All methods are trained on Gibson and evaluated on MP3D. The two settings, “agent-matched goal” and “user-matched goal”, represent the goal camera reflecting the task executor and task publisher, respectively.
@@
Results on Gibson.   For the “agent-matched goal” setting, as shown in Table I, our RSRNav outperforms the state-of-the-art methods, i.e., ranks first for metrics SPL, SR. Our baseline (ZER) achieves 29.2% success, with our RSRNav outperforming it 3-fold. Meanwhile, it achieves 21.6% SPL, and our RSRNav outperforms it by 3.2x. The significant improvement demonstrates the effectiveness of our correlation-based method in ImageNav. Furthermore, our RSRNav shows considerable performance improvements compared to OVRL-V2, which uses a strong visual backbone (ViT) with pre-training. Although employing a simpler backbone, ResNet9, without pre-training, we still perform better than it in both SPL (69.5% vs. 58.7%) and SR (91.1% vs. 82.0%). This suggests that correlation information is more critical for predicting navigation actions than the rich semantic features extracted by complex encoders. Compared to Fgprompt, both using ResNet9, we surpass it by 3.0 for SPL and 0.7 for SR, the improvement in SPL demonstrates that our explicit modeling of correlation and the introduction of directional information can make navigation more efficient.
@@
Cross-Domain Generalization.   Beyond the visual domain disparity among these datasets, MP3D scenes are both more intricate and expansive compared to Gibson, while HM3D showcases a wide variety of scene types. This presents a very challenging cross-domain evaluation setting. Following the setting and testing episodes in ZER, we directly test our Gibson-trained model on these two new datasets without any tuning. As shown in Table II, our RSRNav achieves the best results on SPL and SR under both two goal camera settings in MP3D. It is worth noting that our advantage is amplified under the “user-matched goal” setting, which again proves that our method is more generalizable and robust compared to previous methods. Similar results appear in HM3D, as shown in Table III, although our success rate is lower than FGPrompt for setting “agent-matched goal” by 1.3, it is significantly higher in setting “user-matched goal” by 6.9.
TABLE IV: Comparison of our three progressive correlation-based ImageNav methods on Gibson. Baseline* represents the feature map version baseline. All ablation studies are training 100 million steps using the Gibson dataset. “MC” represents the Minimalist Correlation. “DC” represents the Dense Correlation. “Direct-C” represens the Direction-aware Correlation. All methods are trained on Gibson and evaluated on MP3D. The two settings, “agent-matched goal” and “user-matched goal”, represent the goal camera reflecting the task executor and task publisher, respectively. We report the ImageNav results averaged over three random seeds.
@@
TABLE V: Ablation studies of “Dense Correlation” on Gibson.
@@
TABLE VI: Ablation studies of “Direction-aware Correlation” on Gibson.
@@
IV-C Ablation Study
The “Navigation Via Minimalist Correlation” version is designed to explore the importance of correlation among a large pile of information, which only passes two correlation scores to the policy network. As illustrated in Table IV, it outperforms Baseline by (+2.7 SPL, +6.4 SR), indicating that despite the low information content of “Minimalist Correlation”, it is more useful and effective and that there is information redundancy in Baseline. With this exciting phenomenon, we further enhance the construction and utilization of correlation and evolve it to the “Navigation Via Dense Correlation” version. Both metrics improved by more than 3x with the richer amount of correlation information, further confirming the critical role of correlation in ImageNav. Also, with the same processing of the feature maps we bring 4× improvements in the success rate and 3× improvements in SPL comparing with Baseline*. Finally, we introduce “Direction-aware Correlation” to provide more direction information to cue ImageNav, which brings a great boost (+8.0 SPL, +6.8 SR).